OpenStudy (anonymous):

What are the next three terms of the pattern 18, 14, 10, 6, ...? -10, -6, -2 2, -1, -5 2, -2, -6 -10, -14, -18

OpenStudy (anonymous):

I will give you a mdeal for the answer!!!

OpenStudy (anonymous):

In each term you are subtracting by 4 You can find this from 18-14 = 4 14-10 = 4 10-6 = 4 Now to the next three terms: 6-4 = 2 2-4 = -2 -2 -4 = -6 Does that make sense?

OpenStudy (anonymous):

2,-2,-6